OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally eighteen cubic meters and is then resized using a scale factor of 2 on each side length?

OpenStudy (anonymous):

Volume of Rectangular Prism:\[V=Bh=lwh\]let \(V_1=18~m^3\) therefore \[18=l_1w_1h_1\]for new volume \(V_2\), it would be\[V_2=l_2w_2h_2\]where \(l_2=2l_1\), \(w_2=2w_1\) and \(h_2=2h_1\) therefore\[V_2=2l_1\times 2w_1\times 2h_1=8\times l_1w_1h_1=8\times18\]\[V_2=144~m^3\]